OpenStudy (anonymous):

You invest $2,000 in an account that has an annual interest rate of 3.2%, compounded annually. How much money will be in the account after 12 years? Round your answer to the nearest whole number.

OpenStudy (cakemeister):

Does this use the formula A=P(1+r)/nt ?

OpenStudy (kl0723):

yep

OpenStudy (kl0723):

\[A = P(1+\frac{ r }{ n })^{nt}\] P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year

OpenStudy (kl0723):

your problem is compounded so it is actually the formula above :)

OpenStudy (kl0723):

@onelove26 let me know if you need more help

OpenStudy (anonymous):

(where t = 1, because it's compounded annually = once a year)

OpenStudy (kl0723):

@mg_omg yes :)

OpenStudy (cwrw238):

wheer n = 1 , t = 15

OpenStudy (cwrw238):

sorry t = 12

OpenStudy (anonymous):

so, @kl0723 the equation would be: A=2,000(1+3.2/12)^1*12 Is that correct? or am I wrong?

OpenStudy (kl0723):

is right but remeber the rate has to be in decimal form, you have it in percentage

OpenStudy (kl0723):

to get the percentage in decimal form just divide the value by 100

OpenStudy (anonymous):

which then I will get: .032

OpenStudy (anonymous):

So how will I exactly go about solving this?

OpenStudy (kl0723):

that is 0.032

OpenStudy (anonymous):

okay. but how will I exactly go about solving this type of equation? @kl0723

OpenStudy (kl0723):

just solve for A, you have all the values you need to do so

OpenStudy (anonymous):

should I divide .032 by 12 first??

OpenStudy (anonymous):

@kl0723